# Network-Realized Model Predictive Control Part II: Distributed Constraint Management

### v1

@@ 6 Distributed Implementation of the MPC Layer
In this section, we show how to ensure global theoretical guarantees for the first layer’s closed-loop system, by independently designing an MPC controller each area’s dynamics described in (16a)-(16d). The results presented here are inspired by the core idea investigated in [6], and our proposed technique represents the MPC-oriented analogue of the indicated work.
6.1 Formulating Separate Optimization Problems
In order to compute the second layer’s command signal vector in a distributed manner, we will assign to each of the network areas a set of standard MPC-like control laws, described by the optimization problems from (42), located at the top of this page, where we have that:
  1. a)
    the real-valued weights denoted $c_{i}[t]\in\mathbb{R}^{2n_{xi}+n_{ui}}$, $Q_{i}[t]=Q_{i}[t]^{\top}\in\mathbb{R}^{(2n_{xi}+n_{ui})\times(2n_{xi}+n_{ui})}$, $r_{i}[t]\in\mathbb{R}$ are tuning parameters of the stage costs from (42);
  2. b)
    the scalars $\0T_{i}\in\mathbb{N}\backslash\{0\}$ and $\overline{T}_{i}\in\mathbb{N}$ are, respectively, the (set-)constrained and the unconstrained prediction horizons of each area’s MPC controller;
  3. c)
    the sets ${\mathcal{U}}_{s1i}$ and ${\mathcal{U}}_{s2i}$ are those from (30a)-(30b);
  4. d)
    the sets $\Xi_{it}$ are those defined in (43), located at the top of this page, for any matrices $H_{1it}\in\mathbb{R}^{q_{1it}\times n_{xi}}$, any vectors $H_{2\ell t}\in\mathbb{R}^{q_{2\ell t}}$ and $h_{1it}\in\mathbb{R}^{q_{1it}}$, and any scalars $h_{2\ell t}\in\mathbb{R}$ which satisfy
    $\left\{\begin{split}&\mathcal{P}_{1it}:=\{v\in\mathbb{R}^{n_{xi}}\ |\ H_{1it}v\leq h_{1it}\},\\ &\mathcal{P}_{1it}\subseteq\widetilde{\mathcal{X}}_{it}\ominus(\Psi_{xit}\oplus\widetilde{\Delta}_{xit}),\end{split}\right.\vspace{-3mm}$ (44)
@@ 6.1 Formulating Separate Optimization Problems
    $\left\{\begin{split}&\mathcal{P}_{2\ell t}:=\{v\in\mathbb{R}\ |\ H_{2\ell t}v\leq h_{2\ell t}\},\\ &\mathcal{P}_{2\ell t}\subseteq\widetilde{W}_{\ell 1t}\ominus\big{(}\big{(}e^{\top}_{(\ell-\alpha_{ui})}\Psi_{uit}\big{)}\oplus\big{(}e^{\top}_{(\ell-\alpha_{ui})}\widetilde{\Delta}_{uit}\big{)}\big{)},\end{split}\right.$ (45)
    for every $\ell\in\mathcal{A}_{ui}$.
@@
We implicitly assume that the first-layer closed-loop system has been designed in such a way as to ensure that the sets from (43)-(45) are non-empty. In this scenario, we point out that the MPC formulations provided in (42) are particularly flexible. The tuning parameters $c_{i}[t]$ and $r_{i}[t]$ mentioned in point a) enable the explicit inclusion of reference signals (for the predicted values of an area’s state variables) within the state costs of (42). Indeed, it is precisely due to the time-varying nature of such references that the aforementioned parameters are denoted as being explicit functions of the time index $t$. Moreover, the two prediction horizons from point $a)$ permit the careful management of the computational burdens associated with solving the problems from (42). Since we are interested in satisfying the latter’s constraints only for the sake of recursive feasibility, it is possible to improve closed-loop performance by increasing the unconstrained prediction horizon without also dramatically impacting the overall computational cost.
Each of these MPC subcontrollers will be responsible for independently regulating the dynamics of the area they are assigned to, with the aim of satisfying the constraints located on the right-hand side of (26). Assuming that every problem described in (42) is feasible for every time instant $k\geq k_{0}$ (a fact which will be rigorously addressed in the sequel), each subcontroller can be implemented according to Algorithm 1, located on the next page.
$M_{i\rho_{i}}:=\tiny\begin{bmatrix}F_{i0}&O&O&\dots&O\\ F_{i1}&F_{i0}&O&\dots&O\\ F_{i2}&F_{i1}&F_{i0}&\dots&O\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ F_{i(\rho_{i}-1)}&F_{i(\rho_{i}-2)}&F_{i(\rho_{i}-3)}&\dots&F_{i0}\\ \end{bmatrix}\normalsize,\ F_{ik}:=\begin{bmatrix}C_{xi}\\ C_{ui}\end{bmatrix}\normalsize A_{si}^{k}\begin{bmatrix}B_{s1i}&B_{s2i}\end{bmatrix}\normalsize,\,\forall\,k\in\{0:\rho_{i}-1\}.$ (46)
@@
It is straightforward to notice that the matrix pairs $(H_{1it},h_{1it})$ and $(H_{2\ell t},h_{2\ell t})$ represent polyhedral inner-approximations of the sets on the right-hand sides of (44)-(45). Crucially, these pairs can always be determined offline, which greatly simplifies the online construction of the sets from (43) in Step 3 of Algorithm 1. Moreover, we point out that this feature further reinforces the computational benefits of our architecture.
Although the MPC subcontrollers presented in this section repeatedly interact with their counterparts from neighbouring areas, with the aim of exchanging local state measurements, we highlight the fact that all of the individual command signals are computed separately, in a distributed manner. In addition to this feature, the following result shows that the local actuation efforts of these subcontrollers ensure global theoretical guarantees, with respect to the network’s full dynamics in (4a).
Data: The $i^{\text{th}}$ network’s area model and the design parameters of the optimization problem from (42)
Result: The values of $u_{si1}[k_{0}]$ and $u_{si2}[k_{0}]$
Step 1b: Receive the noise-affected initial condition vectors $\begin{bmatrix}\widetilde{x}_{cj}^{\top}&\widetilde{w}_{cj}^{\top}\end{bmatrix}^{\top}$ from the subcontrollers of all network areas of index $j\in\mathcal{N}_{i}\backslash\{i\}$;
Step 2: Form the vectors $\widetilde{\theta}_{xi}[k_{0}+t]$ and $\widetilde{\theta}_{ui}[k_{0}+t]$ for all $t\in\{1:T_{i}\}$ and use them to construct the sets in (43);
Step 3: Solve (42) for two sequences of optimal command signal vectors $u^{\star}_{s1i}[k_{0}+t-1]$ and $u^{\star}_{s2i}[k_{0}+t-1]$, in which $t\in\{1:T_{i}+\overline{T}_{i}\}$;
Step 4a: Set $u_{s1i}[k_{0}]\leftarrow u^{\star}_{s1i}[k_{0}]$ and $u_{s2i}[k_{0}]\leftarrow u^{\star}_{s2i}[k_{0}]$, then transmit $u_{s1i}[k_{0}]$ to the subcontrollers of all network areas of index $j\neq i$, for which $i\in\mathcal{N}_{j}$;
Step 4b: Receive $u_{s1j}[k_{0}]$ from the subcontrollers of all network areas of index $j\in\mathcal{N}_{i}\backslash\{i\}$, then pass them along to the local first layer subcontroller;
Step 5: Wait for the start of the next sampling period and return to Step 1;
Algorithm 1 Implementation of the $i^{\text{th}}$ subcontroller
@@ 2. A2)
    All of the optimization problems described in (42) are feasible at time $k=k_{0}$.
@@
We point out the fact that Assumption A1) from Proposition 6.1 is in no way restrictive, from a practical standpoint. Indeed, the state variables of the NRF-based layer’s distributed implementations can always be initialized at will, during the control architecture’s startup. Moreover, if one wishes to maintain the network’s state variables in the neighbourhood of a chosen operating point, then it must be possible to first steer the dynamics to that operating point.
When starting from a point inside the state-constraining set $\mathcal{X}$, Proposition 6.1 shows that our control scheme is able to maintain the first layer’s state variables inside $\mathcal{X}$ while also satisfying the command constraints encoded by $\mathcal{U}$, provided that all of the MPC-based problems from (42) are feasible. Given the critical importance of satisfying Assumption A2) from Proposition 6.1, we now proceed to investigate a set of conditions which offer strong theoretical guarantees regarding the feasibility of the optimization problems from (42).
@@ 6.2 Guaranteeing Recursive Feasibility
In order to extend the theoretical guarantees obtained in Proposition 6.1 beyond the initial time instant, we must ensure that each of the MPC-based control laws from (42) are feasible for every time instant $k\in\mathbb{N}$ with $k\geq k_{0}$. This property, known as recursive feasibility (see Section 12.3 in [4]), plays a key role in our chosen approach and, consequently, the following theorem (which directly builds upon the results obtained in Proposition 6.1) represents the paper’s main theoretical result.
@@
For each of the area dynamics described in (16a)-(16d), let an MPC subcontroller be implemented as in Algorithm 1 and let the communicated information between these subcontrollers be affected as in (20). For all $i\in\{1:N\}$, let $\rho_{i}\in\mathbb{N}$ with $\rho_{i}\geq 1$ and define the block-Toeplitz matrices from (46), located at the top of this page, along with the following sets
@@
$\widehat{\Theta}_{i\rho_{i}}:=$$\left\{\begin{bmatrix}\theta_{1}^{\top}&\dots&\theta_{\rho_{i}}^{\top}\end{bmatrix}^{\top}\Big{|}\ \theta_{t}\in\widetilde{\Theta}_{it},\,t\in\{1:\rho_{i}\}\right\},$ (47b)
$\widehat{\mathcal{U}}_{i\rho_{i}}:=$$\Big{\{}\begin{bmatrix}u_{1}^{\top}&\dots&u_{(2\rho_{i})}^{\top}\end{bmatrix}^{\top}\Big{|}\ u_{(2t-1)}\in\mathcal{U}_{s1i},$
$\qquad\qquad\qquad u_{(2t)}\in\mathcal{U}_{s2i},\,t\in\{1:\rho_{i}\}\Big{\}},$ (47c)
@@
$\qquad\quad\left.p_{uj}\in{\mathcal{P}}_{2(j+\alpha_{ui})t},\,j\in\{1:n_{ui}\}\right\},$ (47d)
$\widehat{\mathcal{P}}_{i\rho_{i}}:=$$\left\{\begin{bmatrix}p_{1}^{\top}&\dots&p_{\rho_{i}}^{\top}\end{bmatrix}^{\top}\Big{|}\ p_{t}\in\widetilde{\mathcal{P}}_{it},\,t\in\{1:\rho_{i}\}\right\}.$ (47e)
@@ 2. A2)
    $\widehat{\Theta}_{i\rho_{i}}\subseteq\left((-M_{i\rho_{i}})\,\widehat{\mathcal{U}}_{i\rho_{i}}\right)\oplus\widehat{\mathcal{P}}_{i\rho_{i}}.\vspace{-3mm}$ (48)
@@ 1. $i)$
    Each of the optimization problems described in (42) becomes recursively feasible;
@@ 2. $ii)$
Note that Theorem 6.1 can be readily seen as an MPC-oriented analogue of the main result from [6]. The $N$ areas which are dynamically decoupled by the first layer may be assimilated with the clusters depicted in Figure 2 of [6], whereas the MPC-based subcontrollers of our second layer act as the supervisory controllers depicted in the aforementioned figure. In both cases, global guarantees are achieved when each of the high-level control laws satisfies local and mutually independent conditions.
In order to summarize the design workflow of our control architecture’s second layer, we provide the following high-level algorithm, which condenses the various topics discussed in this paper into a sequence of steps.
@@
Initialization: The state-space dynamics (16a)-(16d);
@@
Step 3c: Compute the polyhedral inner-approximations given in (44)-(45) and the sets from (47a)-(47e);
@@
Before moving on, we reiterate the fact that all of the steps presented in Algorithm 2 (along with the design phase of the first layer, see [10]) can be performed offline, and that any reconfiguration of the second layer’s constraints or cost functions does not impact the first layer’s control laws (in contrast to [1]-[2]). Moreover, due to the separate implementations described in (8) and in Algorithm 1, the runtime cost of our architecture is divided among the network’s areas, which greatly reduces the computational burden of local implementation. Finally, we point out that the aforementioned burden is appealingly modest, since it boils down to several inexpensive matrix-vector multiplications and vector-vector additions, needed to implement (8), along with the solving of (42). Note that the latter is a standard Quadratic Programming optimization problem whenever the sets from (30a)-(30b) are chosen to be polyhedral in nature.

### v4

@@ 6 Distributed Implementation of the MPC Layer
In this section, we proceed to show how to ensure global theoretical guarantees for the first layer’s closed-loop system, by independently implementing an MPC controller for each area’s dynamics described in (16a)-(16d). The results presented here are inspired by the core idea investigated in [AI], and our proposed technique represents the MPC-oriented analogue of the indicated work.
Remark 6.1.
As discussed in this manuscript’s companion paper [part1], our approach showcases a series of notable similarities with the SLS-based procedure from [DMPC1, DMPC2]. Indeed, in both instances, the focus rests upon designing the closed-loop response of the network in accordance with a number of structural restrictions. Additionally, the d-locality constraints from [DMPC1, DMPC2] closely match our own area-based sparsity patterns discussed in Remark 4.4, while the feasibilities of these two approaches are also closely intertwined, since both the SLS [SLS] and the Youla Parametrisation (upon which the first layer is based, see [NRF]) yield the set of all achievable and stable closed-loop maps (see [Luca3]). On the other hand, our procedure leverages the fact that the NRF-based layer can always be designed and implemented offline, thus relegating all online computation to the significantly less resource-demanding MPC layer, as shown in the sequel. This stands in stark contrast to the approach from [DMPC1, DMPC2], where sparsity constraints must be factored into the online computation, alongside all the MPC-like constraints.
6.1 Distributed Enforcement of Constraint Satisfaction
We begin by addressing the final and most important component of the formulations introduced in (8). Recall that, in stating the objectives of our control architecture’s second layer, the optimisation problems from (8) make use of the sets $\Xi_{it}$ to ensure the satisfaction of the constraints pertaining to $x[k]$ and $u[k]$. The aforementioned sets, whose precise expressions have been deferred up to this point, may now finally be stated in a formal manner. This is done by first computing the sets
[…]
@@ 6.1 Distributed Enforcement of Constraint Satisfaction
[…]
for every index $\ell\in\mathcal{A}_{ui}$ and any set of real-valued tuples $\scriptsize\begin{array}[]{c}(H_{1it},h_{1it})\in\mathbb{R}^{q_{1it}\times n_{xi}}\times\mathbb{R}^{q_{1it}}\end{array}$ and $\scriptsize\begin{array}[]{c}(H_{2\ell t},h_{2\ell t})\in\mathbb{R}^{q_{2\ell t}}\times\mathbb{R}\end{array}$ which satisfy (42)-(43). These tuples allow us to define $\Xi_{it}$ as in (44), located at the top of this page.
@@
We implicitly assume that the first-layer closed-loop system has been designed in such a way as to ensure that the sets from (42)-(44) are non-empty. If this is not the case, then the reason is the fact that the first layer has not been appropriately tuned in order to tackle cross-coupling and the exogenous disturbance. Provided that the sets $\mathcal{X}$ and $\mathcal{U}$ (along with their area-based decompositions) are available during the first layer’s design phase, a judicious choice of systemic norm optimisation in Section 5.2 of [part1] (tied to the nature of these sets) can be used to ensure non-emptiness. Indeed, given that our first layer is specifically designed to improve upon the open-loop response characteristics of the network described by (1a)-(1c), our assumption is by no means restrictive. Moreover, it highlights one of the key benefits of our approach: the fact that constraint feasibility is certified after first-layer synthesis, with any potential infeasibility being detected early on during the design procedure.
With all of their components now clarified, and assuming that every problem described in (8) is feasible [^3] This property will be rigorously addressed in the sequel. for each time instant $k\geq k_{0}$, each MPC-based subcontroller in our second layer can be implemented as per Algorithm 1.
Data: The $i^{\text{th}}$ network’s area model and the design parameters of the optimisation problem from (8)
Result: The values of $u_{si1}[k]$ and $u_{si2}[k]$
Step 1b: Receive the noise-affected initial condition vectors $\begin{bmatrix}\widetilde{x}_{cj}^{\top}&\widetilde{w}_{cj}^{\top}\end{bmatrix}^{\top}$ from the subcontrollers of all network areas of index $j\in\mathcal{N}_{i}\setminus\{i\}$;
Step 2: Form the vectors $\widetilde{\theta}_{xi}[k+t]$ and $\widetilde{\theta}_{ui}[k+t]$ for all $t\in\{1:T_{i}\}$ and use them to construct the sets in (44);
Step 3: Solve (8) for two sequences of optimal command signal vectors $u^{\star}_{s1i}[k+t-1]$ and $u^{\star}_{s2i}[k+t-1]$, in which $t\in\{1:T_{i}+\overline{T}_{i}\}$;
Step 4a: Set $u_{s1i}[k]\leftarrow u^{\star}_{s1i}[k]$ and $u_{s2i}[k]\leftarrow u^{\star}_{s2i}[k]$, then transmit $u_{s1i}[k]$ to the subcontrollers of all network areas of index $j\neq i$, for which $i\in\mathcal{N}_{j}$;
Step 4b: Receive $u_{s1j}[k]$ from the subcontrollers of all network areas of index $j\in\mathcal{N}_{i}\setminus\{i\}$, then pass them along to the local first-layer subcontroller;
Step 5: Wait for the start of the following sampling period and return to Step 1;
Algorithm 1 Implementation of the $i^{\text{th}}$ subcontroller
@@
It is straightforward to notice that the tuples formed by $(H_{1it},h_{1it})$ and $(H_{2\ell t},h_{2\ell t})$ characterise polyhedral inner-approximations of the sets on the right-hand sides of (42)-(43). Crucially, these pairs can always be determined offline, which greatly simplifies the online construction of the sets from (44) in Step 3 of Algorithm 1. In addition to this computational benefit, note that no explicit construction of any invariant sets occurs in our setting. For especially large or highly interconnected networks, the type of invariant sets discussed in [R8A] or [R8C] may be highly complex, or even numerically intractable for implementation. Our framework allows us to bypass explicit invariance-related concerns, and it allows us to control the complexity of the sets in (44), by means of the inner approximations given in (42)-(43).
Although the implementations proposed in Algorithm 1 show that the MPC-based subcontrollers interact with their counterparts from neighbouring areas, with the aim of exchanging local state measurements, we highlight the fact that all of the individual command signals are computed separately, in a distributed manner.
Remark 6.4.
Regarding the computation of said command signals, the MPC formulations provided in (8) are particularly flexible, in the context of (distributed) constraint satisfaction. The two prediction horizons, introduced in point c) of Section 3.2 and associated with these problems, permit the careful management of the computational burdens which come with solving (8). Since we are chiefly interested in satisfying the latter’s constraints only for the sake of recursive feasibility (see the numerical example presented in the sequel), it is possible to improve closed-loop performance of the MPC-based layer by increasing the unconstrained prediction horizon without dramatically impacting the overall computational cost.
The following result shows that the local efforts of these subcontrollers ensure global theoretical guarantees.
@@ 2. A2)
    All of the optimisation problems described in (8) are feasible at time $k=k_{0}$.
@@
Note that Assumption A1) from Proposition 6.1 hinges upon the initial condition of the controlled network and upon the sets $\mathcal{X}_{i}$ in (26), since the initial conditions of the implementations in (10) can be imposed at will. Should the system from (1a)-(1c) be initialised in some $x_{nf}\not\in\mathcal{X}$, the procedure presented in this paper and in [part1] can also be used as a recovery mechanism. The key is to relax the state constraints given by $\mathcal{X}$ into some feasible and connected set $\mathcal{X}_{r}\supseteq(\mathcal{X}\cup\{x_{nf}\})$, while keeping the same $\mathcal{U}$ and resynthesising the second layer in accordance with the relaxed constraints. Since, in our framework, the first layer is the one responsible for disturbance rejection and reference tracking, the NRF-based control laws can then be used to drive the network’s state within $\mathcal{X}$, all while the second layer handles command limitation. The first layer’s subcontrollers can then be reinitialised to satisfy Assumption A1), and nominal operation may commence using the original second layer.
When starting from a point inside the state-constraining set $\mathcal{X}$, Proposition 6.1 shows that our control scheme is able to maintain the first layer’s state variables inside $\mathcal{X}$ while also satisfying the command constraints encoded by $\mathcal{U}$, provided that all of the MPC-based problems from (8) are feasible. Given the critical importance of satisfying Assumption A2) from Proposition 6.1, we now proceed to investigate a set of conditions which offer strong theoretical guarantees regarding the feasibility of the optimisation problems from (8).
@@ 6.2 Guaranteeing Recursive Feasibility
In order to extend the theoretical guarantees obtained in Proposition 6.1 beyond the initial time instant, we must ensure that each of the MPC-based control laws from (8) is feasible for every time instant $k\in\mathbb{N}$ with $k\geq k_{0}$. This property, known as recursive feasibility (see Section 12.3 in [BBM]), plays a key role in our chosen approach and, consequently, the following theorem (which directly builds upon the results obtained in Proposition 6.1) represents the paper’s main theoretical result.
@@
For each of the area dynamics described in (16a)-(16d), let an MPC subcontroller be implemented as in Algorithm 1 and let the communicated information between these subcontrollers be affected as in (20). For all $i\in\{1:N\}$, let $\rho_{i}\in\mathbb{N}$ with $\rho_{i}\geq 1$ and define
@@
$\qquad\quad\left.p_{uj}\in{\mathcal{P}}_{2(j+\alpha_{ui})t},\,j\in\{1:n_{ui}\}\right\}.$ (45b)
@@ 2. A2)
    $\hskip-28.45274pt\scriptsize\begin{array}[]{l}\widetilde{\Theta}_{it}\subseteq\hskip-1.42262pt\left(\hskip-1.42262pt\begin{bmatrix}-C_{xi}\\ -C_{ui}\end{bmatrix}\displaystyle\bigoplus_{j=1}^{t}A_{si}^{j-1}\Big(\hskip-1.42262pt(B_{s1i}\hskip 1.42262pt\mathcal{U}_{s1i})\hskip-1.42262pt\oplus\hskip-1.42262pt(B_{s2i}\hskip 1.42262pt\mathcal{U}_{s2i})\hskip-1.42262pt\Big)\hskip-1.42262pt\right)\hskip-1.42262pt\oplus\widetilde{\mathcal{P}}_{it},\end{array}\normalsize\hskip-5.69054pt\vskip-5.69054pt$ (46)
    for all $t\in\{1:\rho_{i}\}$.
@@
    Each of the optimisation problems described in (8) becomes recursively feasible;
@@
Notice the fact that it is possible to retrieve the above-mentioned guarantees by taking $\rho_{i}=1$, for every $i\in\{1:N\}$, resulting in especially efficient implementations (see the numerical example from the sequel) from a computational standpoint. Indeed, as the values of $\rho_{i}$ increase, not only does (46) become more difficult to satisfy, but the additional constraints associated with taking $T_{i}=\rho_{i}$ may dramatically increase the cost of solving (8), depending on the complexity of the sets in (44). As discussed in Remark 6.4, taking larger values for $\overline{T}_{i}$ may yield similar degrees of performance as increasing $T_{i}$, without greatly impacting overall computational costs.
To summarise the design workflow of our architecture’s second layer, we provide a high-level algorithm (located at the top of the next page), which condenses the various topics discussed in this paper into a sequence of steps.
Remark 6.7.
Performing Steps 3a-3c in Algorithm 2 makes the paramount importance of our architecture’s first layer readily apparent. It is precisely due to the proper tuning of the NRF-based control laws that area cross-coupling may be diminished (and potentially rendered negligible, for all practical purposes) to such a degree that the MPC-based subcontrollers may be tuned individually via (30a)-(30b). While this effectively bypasses the need to carefully construct the sets from (30a)-(30b) and it crucially prevents the design procedure of the MPC-based subcontrollers from being co-dependent in any significant way, we point out that said procedure does indeed require information pertaining to the first-layer response of other areas. However, this information is fully available once the NRF-based control laws have been obtained, and it is precisely the role of these distributed controllers to ensure the feasibility (recall Remark 6.2) and the relaxed co-dependency of the second layer’s design.
@@
Initialisation: The state-space dynamics (16a)-(16d);
@@
Step 3c: Compute the polyhedral inner-approximations given in (42)-(43) and the sets from (45a)-(45b);
@@
6.3 Computational Considerations
The following discussion is centred around the three core concepts which underpin the computational burden of our procedure, namely: synthesis costs, design scalability, and implementation complexity.
6.3.1 Synthesis Costs
One of the main features associated with Algorithm 2 is the fact that all of its constituent steps (along with the design phase of the first layer, see our companion paper [part1]) can be performed offline, and that any reconfiguration of the second layer’s constraints or cost functions does not impact the first layer’s control laws (in contrast to the technique from [DMPC1, DMPC2], recalling Remark 6.1). The chief computational burden associated with our solution is represented by the cost of computing the sets from (34a)-(34b) along with (42)-(43) for large-scale applications. Although the measures described in Remarks 5.5 and 6.3 can alleviate some of the complexity associated with these operations, the latter remain heavily dependent on the shape and the structure of the involved sets, making it challenging to come up with a one-size-fits-all computational cost. In spite of this, switching to a homothetic [homot] or even zonotope-based [zonot] framework (while employing various approximations, either in terms of the operation or of the operands themselves) is guaranteed to have a significant, positive impact on the complexity of the synthesis procedure. For a practical and specific example of the costs associated with the design of the second layer, see the numerical example presented in the sequel, and the source files from its GitHub repository.
6.3.2 Design Scalability
While all of the steps which form Algorithm 2 can, indeed, be iterated in a fully offline setting, yet another appealing feature is the fact that once the second layer’s command budgeting has been negotiated via the selection of the sets from (30a)-(30b), the subsequent set-based computations from Algorithm 2 can be performed in parallel. As indicated in Remark 6.7, the design procedure of each area can proceed independently of the others, once the aforementioned choice has been made. This fact is a direct consequence of not employing any notions of joint set invariance (see, for example, [R8C]) and of relying, instead, upon the decoupling-focused control laws discussed in our companion paper [part1].
6.3.3 Implementation Complexity
Due to the separate implementations in (10) and in Algorithm 1, a crucial observation is the fact that the runtime cost of our architecture is divided among the network’s areas. Said burden is appealingly modest, since it reduces to several inexpensive operations, namely: four matrix-vector multiplications and two vector-vector additions for an area’s first-layer subcontroller (10), along with the solving of (8) for the second layer. We point out that the latter problem boils down to a standard Quadratic Programming (QP) optimisation problem, whenever the cost in (8) has a quadratic (and symmetrical) expression and the sets from (30a)-(30b) are chosen to be polyhedral in nature (which is always possible, as per Remark 5.3). Despite the fact that the overall complexity of this problem is highly dependent upon the number of constraints it includes (recall Remark 6.3), note that numerous solvers and optimisation toolboxes (such as [mosek] and [YALMIP], respectively) are widely available for the efficient solving of such problems.